how do i get an answer for 4y+y+1=7(y-1)?
im having alot of truble with algebra im a sophmore still taking algebra-1

Answer by haileytucki(390)

4y+y+1=7(y-1)
Since 4y and y are like terms, add y to 4y to get 5y.
5y+1=7(y-1)
Multiply 7 by each term inside the parentheses.
5y+1=7y-7
Since 7y contains the variable to solve for, move it to the left-hand side of the equation by subtracting 7y from both sides.
5y+1-7y=-7
Since 5y and -7y are like terms, add -7y to 5y to get -2y.
-2y+1=-7
Since 1 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 1 from both sides.
-2y=-1-7
Subtract 7 from -1 to get -8.
-2y=-8
Divide each term in the equation by -2.
-(2y)/(-2)=-(8)/(-2)
Simplify the left-hand side of the equation by canceling the common factors.
y=-(8)/(-2)
Simplify the right-hand side of the equation by simplifying each term.
y=4